OpenStudy (anonymous):

Write the slope intercept equation for the line containing the given pair of points..(-3,4) and (3,1).... Please HELP ME ANYONE

hero (hero):

1. Find Slope 2. Use Point-Slope Formula 3. Simplify

OpenStudy (anonymous):

I dont know the formula

OpenStudy (anonymous):

slope = m m= (y2-y1)/(x2-x1)

OpenStudy (anonymous):

yep, than plug m in: y-y1= m(x-x1) to find eq.....

OpenStudy (anonymous):

can you show me where each number goes and I will do the rest I get confused with which is the y-y1 and the x-x1.

hero (hero):

Trust, it must still be early where you're at.

OpenStudy (anonymous):

Usa,

OpenStudy (anonymous):

I live in the carolinas

hero (hero):

Then that means it's very early.

hero (hero):

and very late

OpenStudy (anonymous):

like 3am here where you at in the world. and what is the time there

OpenStudy (anonymous):

x1 and y1 is from one of your two given points. (-3,4) = (x1,y1) & (3,1)= (x2,y2). What did u get for slope m?

OpenStudy (anonymous):

yes 1/2 but where do I go next

OpenStudy (anonymous):

You should get -1/2 for m...?

hero (hero):

Late as ever.

OpenStudy (anonymous):

i did what is next

OpenStudy (anonymous):

Now, use eq: y-y1=m(x-x1) to find slope eq.

OpenStudy (anonymous):

You could use either points, both will give u the same answer.

OpenStudy (anonymous):

ok how do i start

OpenStudy (anonymous):

Just plug in y1 and x1 in the eq. and solve using algebra.

OpenStudy (anonymous):

so -3=-1/2(4)+b

OpenStudy (anonymous):

I believe you using the other eq. "y=mx+b" b is your y-intercept on the graph (if u graphed it). You are not trying to find b but instead an equation of a line in that form. Use this my friend: y-y1= m(x-x1) * You know m * You know y1, x1 * Plug and chug.

OpenStudy (anonymous):

so it would be 5=1/2(6)

OpenStudy (anonymous):

3=1/2(6)

OpenStudy (anonymous):

Uhmm, not quite: y-y1= m(x-x1) Lets do the first point: (-3,4) ---> (x1, y1) Plug in what u know: y1= 4, x1=-3, m=-1/2 y-4= (-1/2)(x-(-3)) From here, solve...

OpenStudy (anonymous):

um -4=-1-1... i dont know I hate math. ugh

OpenStudy (anonymous):

will you please write out everything foe me i am already gonna give you the medal please with cherries on top.

OpenStudy (anonymous):

you dont do anything with y and x. at the end, u should have something like y= ?x + ?

OpenStudy (anonymous):

i will simplify u one more step to see if u could finish it by yourself: y-4= (-1/2)(x-(-3)) y-4=-.5x -1.5 easier now? solve

OpenStudy (anonymous):

oh I think I got it now y=1/2x+5/2

OpenStudy (anonymous):

yes cept your slope (m) is -.5. if u plug in the other point, you will have the same answer. your 5/2 is your y intercept. if u wanna see, plug 0 for x.

OpenStudy (anonymous):

o okay is now thank you

OpenStudy (anonymous):

your welcome. bye bye

OpenStudy (anonymous):

bye